Two liters of 40% alcohol are mixed with 4 liters of another alcohol to form a 50 % alcohol solution. what is the strength of the added alcohol?
Answer by fractalier(6550)

Let x be the percentage of the added alcohol.
The setup looks like this:
2(40) + 4(x) = 6(50)
Simplifying and solving we have
80 + 4x = 300
4x = 220
x = 55%